# Badminton Birdie-Like Aerodynamic Alignment of Drifting Dust Grains by Subsonic Gaseous Flows in Protoplanetary Disks

### v1

@@ 3.2 Strictly Axisymmteric Spheroid
$D_{s}$ $=\rho_{g}v_{\text{th}}\pi c^{4}\bar{a}\bigg{[}\bar{g}^{2}E[1-\bar{z}^{2}]+(1-\bar{a}^{2})^{2}E[\bar{z}^{2}-\bar{z}^{4}]\bigg{]}$
$P_{s}$ $=\rho_{g}v_{\text{th}}A\pi c^{3}\bar{a}\bar{g}E[1-\bar{z}^{2}]$
@@
We find that the potential energy is $U_{s}=-P_{s}\cos\theta$, which is again proportional to $-\cos\theta$ and means the final alignment angle is at $\theta=0$ for $P_{s}>0$ or at $\theta=\pi$ for $P_{s}<0$. Since $E[1-\bar{z}^{2}]$ is always positive for all $\bar{a}>0$, the sign of $P_{s}$ comes entirely from the sign of $g$ and is independent of $\bar{a}$. In other words, both prolate and oblate grains will become aligned with their axis of symmetry along the gas flow, such that the direction from the center of mass to the geometric center follows the direction of the gas flow ($\bm{g}\parallel\bm{A}$).
@@
$t_{d,s}$ $=\frac{\rho_{s}c}{\rho_{g}v_{\text{th}}}\hat{t}_{d,s},$
$\hat{t}_{d,s}$ $\equiv\frac{8}{15}\frac{\bar{a}(1+\bar{a}^{2}+5\bar{g}^{2})}{\bar{g}^{2}E[1-\bar{z}^{2}]+(1-\bar{a}^{2})^{2}E[\bar{z}^{2}-\bar{z}^{4}]}.$ (60)
The damping time $t_{d,s}$ is once again related to the stopping time modified by a dimensionless quantity that depends on the structure of the grain $\hat{t}_{d,s}$ (through $\bar{a}$ and $\bar{g}$). Fig. 6a shows how $\hat{t}_{d,s}$ depends on $\bar{a}$ and $\bar{g}$. If we have a sphere ($\bar{a}=1$) rotating about its geometric center ($\bar{g}=0$), $\hat{t}_{d,s}\rightarrow\infty$ (or equivalently, $D_{s}\rightarrow 0$) meaning it will not be damped by the gas at all. The result makes sense since the normal direction at each point of the sphere is always directed radially from the center of the sphere and thus the gas cannot produce any torque as the sphere spins. However, if the center of mass of the sphere does not correspond to the geometrical center ($\bar{g}\neq 0$), then $t_{d,s}$ is no longer infinite and its rotation can be damped. If we have a homogenous spheroid ($\bar{a}\neq 1$) rotating about its geometric center ($\bar{g}=0$), $\hat{t}_{d,s}$ is also finite since the normal directions of most of the surface elements are no longer strictly in the radial direction from $O$ and gas can produce a torque. In general, as the non-sphericity ($|\bar{a}-1|$) increases, $\hat{t}_{d,s}$ decreases. In other words, more non-spherical grains (for both prolates and oblates) are better damped by the gas and are aligned faster. With $\bar{g}=0.01$ and $\bar{a}=0.9$, $\hat{t}_{d,s}\sim 80$.
Figure 6: Dimensionless factors for the strictly axisymmetric spheroid as a function of $\bar{g}$ (the normalized distance from the center of mass to the spheroid center, $g/c$) and $\bar{a}$ (the aspect ratio, $a/c$). Panel a: the damping time dimensionless factor, $\hat{t}_{d,d}$. Panel b: the oscillation time dimensionless factor, $\hat{t}_{o,d}$. Black solid lines correspond to the spherical case, $\bar{a}=1$. Dashed lines correspond to prolates ($\bar{a}<1$), while dash-dotted lines correspond to oblates ($\bar{a}>1$).
@@
$t_{o,s}$ $=\sqrt{\frac{\rho_{s}c^{2}}{\rho_{g}v_{\text{th}}A}}\hat{t}_{o,s}$
$\hat{t}_{o,s}$ $\equiv\sqrt{\frac{16\pi^{2}}{15}\frac{\bar{a}(1+\bar{a}^{2}+5\bar{g}^{2})}{\bar{g}E[1-\bar{z}^{2}]}}$ (61)
which is also similar to the expression for the double-sphere. Fig. 6b shows how $\hat{t}_{o,s}$ depends on the structure. For a given $\bar{a}$, increasing $\bar{g}$ decreases $t_{o,s}$ for both prolates and oblates, which means the key to allowing a spheroid to oscillate lies in the offset between the center of mass from the geometric center. For $\bar{g}\rightarrow 0$, $\hat{t}_{o,s}\rightarrow\infty$ (or equivalently, $P_{s}\rightarrow 0$) meaning the grain will not oscillate without the offset. Fig. 6b also shows that increasing $\bar{a}$ (more oblong) increases $\hat{t}_{o,d}$. However, the large difference is in part due to keeping the length along the axis of symmetry, $c$, fixed (long axis for a prolate, but short axis for an oblate). With $\bar{g}=0.01$ and $\bar{a}=0.9$, $\hat{t}_{o,s}\sim 35$.
@@ 3.3 Quasi-Axisymmetric Spheroid
Following the same notation above, consider $\bm{g}$ that forms an angle $\psi$ from $\bm{b}_{3}$. Fig. 7a shows a schematic of a quasi-axisymmetric oblate. We pick $\bm{b}_{1}$ to be in the same plane as $\bm{g}$ and $\bm{b}_{3}$, which gives
@@
using $\bm{b}_{i}$ as basis vectors. The gas flow remains in the $\bm{e}_{3}$ direction. We can choose $\bm{e}_{1}$ such that $\bm{g}$ lies in the plane formed by $\bm{e}_{1}$ and $\bm{e}_{3}$. $\bm{e}_{2}$ remains parallel to $\bm{b}_{2}$.
Figure 7: Panel a: a schematic of the coordinate system used to describe the quasi-axisymmetric oblate following the same coloring scheme and notation as Fig. 5. Without strictly requiring $\bm{g}$ to be parallel to $\bm{b}_{3}$, it can form an angle $\psi$ from $\bm{b}_{3}$. The green dotted line helps visualize the angle $\theta$ formed by $\bm{b}_{3}$ and $\bm{e}_{3}$. The blue dotted line is in the direction of $\bm{g}$ which helps visualize $\psi$. Panel b: a schematic of a quasi-axisymmetric oblate with $\bm{g}\parallel\bm{b}_{1}$ (or equivalently $\psi=90^{\circ}$) when reaching alignment at $\theta=-90^{\circ}$.
@@
$\tilde{L}_{11}$ $=-\pi c^{4}\bar{a}\bar{g}^{2}E[1-\bar{z}^{2}]\cos^{2}\psi$
$\tilde{L}_{13}=\tilde{L}_{31}$ $=-\pi c^{4}\bar{a}\bar{g}^{2}E[1-\bar{z}^{2}]\cos\psi\sin\psi$
$\tilde{L}_{22}$ $=-\pi c^{4}\bar{a}\bar{g}^{2}\bigg{[}E[1-\bar{z}^{2}]\cos^{2}\psi+2\bar{a}^{2}E[\bar{z}^{2}]\sin^{2}\psi\bigg{]}$
$\tilde{L}_{33}$ $=-\pi c^{4}\bar{a}\bar{g}^{2}E[1-\bar{z}^{2}]\sin^{2}\psi$ (63)
@@
$\tilde{M}_{12}$ $=-\pi c^{3}\bar{a}\bar{g}E[1-\bar{z}^{2}]\cos\psi$
$\tilde{M}_{21}$ $=\pi c^{3}\bar{a}\bar{g}E[1-\bar{z}^{2}]\cos\psi$
$\tilde{M}_{23}$ $=-2\pi c^{3}\bar{a}^{3}\bar{g}E[\bar{z}^{2}]\sin\psi$
$\tilde{M}_{32}$ $=\pi c^{3}\bar{a}\bar{g}E[1-\bar{z}^{2}]\sin\psi$ (64)
@@
$D_{q}$ $=\rho_{g}v_{\text{th}}\pi c^{4}\bar{a}\bigg{[}(1-\bar{a}^{2})^{2}E[\bar{z}^{2}-\bar{z}^{4}]$ $+\bar{g}^{2}E[1-\bar{z}^{2}]\cos^{2}\psi$
$+2\bar{a}^{2}\bar{g}^{2}E[\bar{z}^{2}]\sin^{2}\psi\bigg{]}$
$P_{q}$ $=\rho_{g}v_{\text{th}}A\pi c^{3}\bar{a}\bar{g}E[1-\bar{z}^{2}]\cos\psi$
$P_{c}$ $=2\rho_{g}v_{\text{th}}A\pi c^{3}\bar{a}^{3}\bar{g}E[\bar{z}^{2}]\sin\psi\text{ .}$
@@ 4.2 Alignment timescale
Fig. 11c shows the timescales as a function of radius in the disk. We find that the damping time $t_{d}$ is around 4 days ($\sim 10^{-2}$ year) at $R=1$ au, and reaches a thousand years at $R=100$ au. Also, $t_{d}$ is mostly smaller than the orbital time $t_{K}$ at all radii up to 70 au, meaning grain oscillations are mostly damped within a fraction of an orbit. The increase in $t_{d}$ with increasing radius is due to the decrease in both the midplane density and temperature (Eq. 3.2). Given that $t_{d}$ is comparable to or less than $t_{K}$, we reason that the alignment mechanism can operate in disks even for a grain with a small offset between its center of mass and its geometric center of $g=0.01$.
The oscillation time $t_{o}$ is around a few minutes ($\sim 10^{-5}$ year) at $R=1$ au and increases to an hour ($\sim 10^{-4}$ year) at $R=100$ au. Different from $t_{d}$, the oscillation time $t_{o}$ does not increase strongly with radius, because it is only proportional inversely to the square root of $\rho_{g}$ and in addition, benefits from an increased level of gas flow at larger radii. From the adopted parameters, the St is $\sim 10^{-3}$ at 1 au and increases with increasing radius to $\sim 0.1$ because the midplane density decreases while the grain size is fixed (Fig. 11a). The increasing St allows the grain to better decouple and experience stronger flow speed $A$ (Fig. 11b).
As discussed previously for the disk and grain parameters chosen in this illustrative example, the direction of the flow is predominantly in the radial direction in the simple power-law prescription of a disk (Fig. 11b). More comprehensive calculations of the gas and dust velocity in structured disks are necessary to address the question of the alignment direction.
Figure 11: The resulting power-law profiles of a disk. Panel a: the St. Panel b: the dashed green and orange lines are the absolute values of the radial and azimuthal components of the gas flow, respectively. The black line is the magnitude of the gas flow $A$. Panel c: the solid line is the damping time $t_{d}$, the dashed line is the oscillation time $t_{o}$, and the dotted line is the Keplerian orbital time $t_{K}$.
@@ 5.1 Comparison to other mechanical alignment mechanisms
Our formulation of Eq. 37 does not immediately produce the results of Gold alignment. The grain structure from Gold (1952a) corresponds to a prolate which has $\bm{K}=\bm{0}$ and also the center of mass is the same as the geometric center, making $\textbf{{M}}=\textbf{{0}}$. As a result, the torque from $\bm{A}$ is always $\bm{0}$ regardless of how large $\bm{A}$ is. Therefore, we would expect that spinning grains damped by the gas should eventually stop at an orientation that is entirely determined by the initial conditions (see Eq. 17). One may notice that from Eq. 17 the spin axis will be around $\bm{e}_{2}$, which is defined to be perpendicular to $\bm{A}$, and we may obtain polarization from the Gold-projection effect (second part). However, the polarization is an artificial result because we have assumed that grains can only oscillate around $\bm{e}_{2}$. In 3D, with $\textbf{{M}}=\textbf{{0}}$, the orientation of angular momenta and the resulting final orientation should be entirely random with respect to $\bm{A}$.
@@ 5.2 Does birdie-like alignment work in the ISM or protostellar envelopes?
$A>A_{t,s}$ $\equiv\frac{fv_{\text{th}}}{n_{g}c^{3}}\hat{A}_{t,s}\text{ ,}$
$\hat{A}_{t,s}$ $\equiv\frac{1}{16\bar{a}\bar{g}E[1-\bar{z}^{2}]}$ (80)
where $\hat{A}_{t,s}$ is a dimensionless factor and the physical combination is the characteristic threshold drift speed. For $\bar{a}=0.9$ and $\bar{g}=0.01$, we get $\hat{A}_{t,d}\sim 5$.
@@
We next assess the dense cores of molecular clouds, since dust polarization is routinely measured in dense cores and protostellar envelopes (e.g. Stephens et al., 2013; Maury et al., 2018; Galametz et al., 2018; Le Gouellec et al., 2019; Yen et al., 2020; Pattle et al., 2021; Cox et al., 2022; Le Gouellec et al., 2023; Lin et al., 2024b; Huang et al., 2024). We find that the damping time is

### v2

@@ 3.2 Strictly Axisymmteric Spheroid
$D_{s}$ $=\rho_{g}v_{\text{th}}\pi c^{4}\bar{a}\bigg{[}\bar{g}^{2}E[1-x^{2}]+(1-\bar{a}^{2})^{2}E[x^{2}-x^{4}]\bigg{]}$
$P_{s}$ $=\rho_{g}v_{\text{th}}A\pi c^{3}\bar{a}\bar{g}E[1-x^{2}]$
@@
We find that the potential energy is $U_{s}=-P_{s}\cos\theta$, which is again proportional to $-\cos\theta$ and means the final alignment angle is at $\theta=0$ for $P_{s}>0$ or at $\theta=\pi$ for $P_{s}<0$. Since $E[1-x^{2}]$ is always positive for all $\bar{a}>0$, the sign of $P_{s}$ comes entirely from the sign of $g$ and is independent of $\bar{a}$. In other words, both prolate and oblate grains will become aligned with their axis of symmetry along the gas flow, such that the direction from the center of mass to the geometric center follows the direction of the gas flow ($\bm{g}\parallel\bm{A}$).
@@
$t_{d,s}$ $=\frac{\rho_{s}c}{\rho_{g}v_{\text{th}}}\breve{t}_{d,s},$
$\breve{t}_{d,s}$ $\equiv\frac{8}{15}\frac{\bar{a}(1+\bar{a}^{2}+5\bar{g}^{2})}{\bar{g}^{2}E[1-x^{2}]+(1-\bar{a}^{2})^{2}E[x^{2}-x^{4}]}.$ (60)
The damping time $t_{d,s}$ is once again related to the stopping time modified by a dimensionless quantity that depends on the structure of the grain $\breve{t}_{d,s}$ (through $\bar{a}$ and $\bar{g}$). Fig. 6a shows how $\breve{t}_{d,s}$ depends on $\bar{a}$ and $\bar{g}$. If we have a sphere ($\bar{a}=1$) rotating about its geometric center ($\bar{g}=0$), $\breve{t}_{d,s}\rightarrow\infty$ (or equivalently, $D_{s}\rightarrow 0$) meaning it will not be damped by the gas at all. The result makes sense since the normal direction at each point of the sphere is always directed radially from the center of the sphere and thus the gas cannot produce any torque as the sphere spins. However, if the center of mass of the sphere does not correspond to the geometrical center ($\bar{g}\neq 0$), then $t_{d,s}$ is no longer infinite and its rotation can be damped. If we have a homogenous spheroid ($\bar{a}\neq 1$) rotating about its geometric center ($\bar{g}=0$), $\breve{t}_{d,s}$ is also finite since the normal directions of most of the surface elements are no longer strictly in the radial direction from $O$ and gas can produce a torque. In general, as the non-sphericity ($|\bar{a}-1|$) increases, $\breve{t}_{d,s}$ decreases. In other words, more non-spherical grains (for both prolates and oblates) are better damped by the gas and are aligned faster. With $\bar{g}=0.01$ and $\bar{a}=0.9$, $\breve{t}_{d,s}\sim 80$.
Figure 6: Dimensionless factors for the strictly axisymmetric spheroid as a function of $\bar{g}$ (the normalized distance from the center of mass to the spheroid center, $g/c$) and $\bar{a}$ (the aspect ratio, $a/c$). Panel a: the damping time dimensionless factor, $\breve{t}_{d,d}$. Panel b: the oscillation time dimensionless factor, $\breve{t}_{o,d}$. Black solid lines correspond to the spherical case, $\bar{a}=1$. Dashed lines correspond to prolates ($\bar{a}<1$), while dash-dotted lines correspond to oblates ($\bar{a}>1$).
@@
$t_{o,s}$ $=\sqrt{\frac{\rho_{s}c^{2}}{\rho_{g}v_{\text{th}}A}}\breve{t}_{o,s}$
$\breve{t}_{o,s}$ $\equiv\sqrt{\frac{16\pi^{2}}{15}\frac{\bar{a}(1+\bar{a}^{2}+5\bar{g}^{2})}{\bar{g}E[1-x^{2}]}}$ (61)
which is also similar to the expression for the double-sphere. Fig. 6b shows how $\breve{t}_{o,s}$ depends on the structure. For a given $\bar{a}$, increasing $\bar{g}$ decreases $t_{o,s}$ for both prolates and oblates, which means the key to allowing a spheroid to oscillate lies in the offset between the center of mass from the geometric center. For $\bar{g}\rightarrow 0$, $\breve{t}_{o,s}\rightarrow\infty$ (or equivalently, $P_{s}\rightarrow 0$) meaning the grain will not oscillate without the offset. Fig. 6b also shows that increasing $\bar{a}$ (more oblong) increases $\breve{t}_{o,d}$. However, the large difference is in part due to keeping the length along the axis of symmetry, $c$, fixed (long axis for a prolate, but short axis for an oblate). With $\bar{g}=0.01$ and $\bar{a}=0.9$, $\breve{t}_{o,s}\sim 35$.
@@ 3.3 Quasi-Axisymmetric Spheroid
Following the same notation above, consider $\bm{g}$ that forms an angle $\psi$ from $\bm{\hat{b}}_{3}$. Fig. 7a shows a schematic of a quasi-axisymmetric oblate. We pick $\bm{\hat{b}}_{1}$ to be in the same plane as $\bm{g}$ and $\bm{\hat{b}}_{3}$, which gives
@@
using $\bm{\hat{b}}_{i}$ as basis vectors. The gas flow remains in the $\bm{\hat{e}}_{3}$ direction. We can choose $\bm{\hat{e}}_{1}$ such that $\bm{g}$ lies in the plane formed by $\bm{\hat{e}}_{1}$ and $\bm{\hat{e}}_{3}$. $\bm{\hat{e}}_{2}$ remains parallel to $\bm{\hat{b}}_{2}$.
Figure 7: Panel a: a schematic of the coordinate system used to describe the quasi-axisymmetric oblate following the same coloring scheme and notation as Fig. 5. Without strictly requiring $\bm{g}$ to be parallel to $\bm{\hat{b}}_{3}$, it can form an angle $\psi$ from $\bm{\hat{b}}_{3}$. The green dotted line helps visualize the angle $\theta$ formed by $\bm{\hat{b}}_{3}$ and $\bm{\hat{e}}_{3}$. The blue dotted line is in the direction of $\bm{g}$ which helps visualize $\psi$. Panel b: a schematic of a quasi-axisymmetric oblate with $\bm{g}\parallel\bm{\hat{b}}_{1}$ (or equivalently $\psi=90^{\circ}$) when reaching alignment at $\theta=-90^{\circ}$.
@@
$\tilde{L}_{11}$ $=-\pi c^{4}\bar{a}\bar{g}^{2}E[1-x^{2}]\cos^{2}\psi$
$\tilde{L}_{13}=\tilde{L}_{31}$ $=-\pi c^{4}\bar{a}\bar{g}^{2}E[1-x^{2}]\cos\psi\sin\psi$
$\tilde{L}_{22}$ $=-\pi c^{4}\bar{a}\bar{g}^{2}\bigg{[}E[1-x^{2}]\cos^{2}\psi+2\bar{a}^{2}E[x^{2}]\sin^{2}\psi\bigg{]}$
$\tilde{L}_{33}$ $=-\pi c^{4}\bar{a}\bar{g}^{2}E[1-x^{2}]\sin^{2}\psi$ (63)
@@
$\tilde{M}_{12}$ $=-\pi c^{3}\bar{a}\bar{g}E[1-x^{2}]\cos\psi$
$\tilde{M}_{21}$ $=\pi c^{3}\bar{a}\bar{g}E[1-x^{2}]\cos\psi$
$\tilde{M}_{23}$ $=-2\pi c^{3}\bar{a}^{3}\bar{g}E[x^{2}]\sin\psi$
$\tilde{M}_{32}$ $=\pi c^{3}\bar{a}\bar{g}E[1-x^{2}]\sin\psi$ (64)
@@
$D_{q}$ $=\rho_{g}v_{\text{th}}\pi c^{4}\bar{a}\bigg{[}(1-\bar{a}^{2})^{2}E[x^{2}-x^{4}]$ $+\bar{g}^{2}E[1-x^{2}]\cos^{2}\psi$
$+2\bar{a}^{2}\bar{g}^{2}E[x^{2}]\sin^{2}\psi\bigg{]}$
$P_{q}$ $=\rho_{g}v_{\text{th}}A\pi c^{3}\bar{a}\bar{g}E[1-x^{2}]\cos\psi$
$P_{c}$ $=2\rho_{g}v_{\text{th}}A\pi c^{3}\bar{a}^{3}\bar{g}E[x^{2}]\sin\psi\text{ .}$
@@ 4.2 Alignment timescale
For reference, we estimate the mean free path of the gas by
$\lambda_{\text{mfp}}=\frac{1}{n_{g}\sigma_{\rm H_{2}}}$ (79)
where $\sigma_{\rm H_{2}}=2\times 10^{-15}$ cm2 (Birnstiel et al., 2010). Fig. 11a shows $\lambda_{\text{mfp}}$ as a function of radius in the disk. For our adopted disk parameters, the $\lambda_{\text{mfp}}$ is $\sim 8$ cm at $R=1$ au to $\sim 2\times 10^{5}$ cm at $=100$ au. Thus, the assumed millimeter grain should be well within the Epstein regime, especially in the lower-density regions like in the outer regions of a disk.
Fig. 11d shows the timescales as a function of radius in the disk. We find that the damping time $t_{d}$ is around 4 days ($\sim 10^{-2}$ year) at $R=1$ au, and reaches a thousand years at $R=100$ au. Also, $t_{d}$ is mostly smaller than the orbital time $t_{K}$ at all radii up to 70 au, meaning grain oscillations are mostly damped within a fraction of an orbit. The increase in $t_{d}$ with increasing radius is due to the decrease in both the midplane density and temperature (Eq. 3.2). Given that $t_{d}$ is comparable to or less than $t_{K}$, we reason that the alignment mechanism can operate in disks even for a grain with a small offset between its center of mass and its geometric center of $g=0.01$.
The oscillation time $t_{o}$ is around a few minutes ($\sim 10^{-5}$ year) at $R=1$ au and increases to an hour ($\sim 10^{-4}$ year) at $R=100$ au. Different from $t_{d}$, the oscillation time $t_{o}$ does not increase strongly with radius, because it is only proportional inversely to the square root of $\rho_{g}$ and in addition, benefits from an increased level of gas flow at larger radii. From the adopted parameters, the St is $\sim 10^{-3}$ at 1 au and increases with increasing radius to $\sim 0.1$ because the midplane density decreases while the grain size is fixed (Fig. 11b). The increasing St allows the grain to better decouple and experience stronger flow speed $A$ (Fig. 11c).
As discussed previously for the disk and grain parameters chosen in this illustrative example, the direction of the flow is predominantly in the radial direction in the simple power-law prescription of a disk (Fig. 11c). More comprehensive calculations of the gas and dust velocity in structured disks are necessary to address the question of the alignment direction.
Figure 11: The resulting power-law profiles of a disk. Panel a: the mean free path of the gas. Panel b: the St. Panel c: the dashed green and orange lines are the absolute values of radial and azimuthal components of the gas flow, respectively. The black line is the magnitude of the gas flow $A$. Panel d: the solid line is the damping time $t_{d}$, the dashed line is the oscillation time $t_{o}$, and the dotted line is the Keplerian orbital time $t_{K}$.
@@ 5.1 Comparison to other mechanical alignment mechanisms
Our formulation of Eq. 37 does not immediately produce the results of Gold alignment. The grain structure from Gold (1952a) corresponds to a prolate which has $\bm{K}=\bm{0}$ and also the center of mass is the same as the geometric center, making $\textbf{{M}}=\textbf{{0}}$. As a result, the torque from $\bm{A}$ is always $\bm{0}$ regardless of how large $\bm{A}$ is. Therefore, we would expect that spinning grains damped by the gas should eventually stop at an orientation that is entirely determined by the initial conditions (see Eq. 17). One may notice that from Eq. 17 the spin axis will be around $\bm{\hat{e}}_{2}$, which is defined to be perpendicular to $\bm{A}$, and we may obtain polarization from the Gold-projection effect (second part). However, the polarization is an artificial result because we have assumed that grains can only oscillate around $\bm{\hat{e}}_{2}$. In 3D, with $\textbf{{M}}=\textbf{{0}}$, the orientation of angular momenta and the resulting final orientation should be entirely random with respect to $\bm{A}$.
@@ 5.2 Does birdie-like alignment work in the ISM or protostellar envelopes?
$A>A_{t,s}$ $\equiv\frac{fv_{\text{th}}}{n_{g}c^{3}}\breve{A}_{t,s}\text{ ,}$
$\breve{A}_{t,s}$ $\equiv\frac{1}{16\bar{a}\bar{g}E[1-x^{2}]}$ (81)
where $\breve{A}_{t,s}$ is a dimensionless factor and the physical combination is the characteristic threshold drift speed. For $\bar{a}=0.9$ and $\bar{g}=0.01$, we get $\breve{A}_{t,d}\sim 5$.
@@
We next assess the dense cores of molecular clouds, since dust polarization is routinely measured in dense cores and protostellar envelopes (e.g. Stephens et al., 2013; Maury et al., 2018; Galametz et al., 2018; Le Gouellec et al., 2019; Yen et al., 2020; Pattle et al., 2021; Cox et al., 2022; Le Gouellec et al., 2023; Lin2024ApJ...961..117L; Huang et al., 2024). We find that the damping time is